Waste water from a cement factory contains 0.25 g of Ca2+ and 0.056 g of Mg2+ per 100.0 L of solution. The solution density is 1.001 g/mL. Calculate the Ca+2 and Mg +2 concentrations in ppt, ppm and ppb (by mass)
1
Expert's answer
2021-02-15T02:52:37-0500

The mass of 100L waste water solution is

(100L solution)×"\frac{1ml}{10^{-3}L}" ×"\frac{1.001g}{1ml}" = 1.001×105 g

ppt = "\frac{mass of solute}{mass of solution}" × 1012

Ca2+ ppt = "\frac{0.25}{1.001×10^5}" ×1012 = 2.5×106 ppt

Mg2+ ppt = "\frac{0.056}{1.001×10^5}" × 1012 = 5.6×105 ppt

ppm = "\frac{mass of solute }{mass of solution}" ×106

Ca2+ ppm = "\frac{0.25}{1.001×10^5}" ×106 = 2.5 ppm

Mg2+ ppm = "\frac{0.056}{1.001×10^5}" × 106 = 0.56 ppm

ppb = "\frac{mass of solute }{mass of solution}" ×109

Ca2+ ppb = "\frac{0.25}{1.001×10^5}" ×109 = 2.5 × 103 ppb

Mg2+ ppb = "\frac{0.056 }{1.001×10^5}" ×109 = 5.6×102 ppb